evalute the expression for: x=3, y=2, z=5: z^2-y^2/xz-2y(z-x)
Answer by ewatrrr(24785)


Hi
evalute the expression for: x=3, y=2, z = 5
z^2-y^2/xz-2y(z-x) |substituting the values of x,y and z
(5^2-2^2)/(15-4*2) = 21/7 = 3